jordan is 3 years less than twice the age of his cousin. if their ages total 48 how old is jordan answer
Answer by ikleyn(52781)

.

Let x = age of his cousin.


Then Jordan's age is (2x-3).


The equation for the sum of their ages is


    x + (2x-3) = 48


Simplify and find x


    3x = 48 + 3

    3x = 51

    x = 51/3 = 17.


ANSWER.  Jordan is  2*17-3 = 31 years old.